OpenStudy (paradise):

Solve the following system of equations by using the substitution method. y = -3x + 5 5x – 4y = -3

hartnn (hartnn):

wherever you see 'y' in 5x – 4y = -3, just replace(substitute) it by -3x+5 because y equals -3x+5 what u get ?

OpenStudy (nirmalnema):

put the value of in the 2nd eq 5x-4(-3x+5)=-3 5x+12x-20=-3 17x=17 x = 1 and y =2

OpenStudy (paradise):

@nirmalnema can you help me with another one?

OpenStudy (nirmalnema):

@Paradise yes sure i will try my best

OpenStudy (paradise):

Okay